OpenStudy (anonymous):

wo people decided to start a business of renting out a reception hall. It cost $19,200 to start up this business. They earn $1200 for each event they rent their hall to. Write the linear function that would be used to determine how many rentals they would need to book in order to break even on their investment.

OpenStudy (anonymous):

y = 1,200x - 19,200 -> x = 19,200/ 1,200 = 16 events to break even.